OpenStudy (anonymous):

For the fundamental solution set S={e^x, e^(2x), e^(3x)} can we construct a linear ODE with constant coefficients?

OpenStudy (anonymous):

I have verified that the set if linearly independent. But after that I am not sure how to proceed.

OpenStudy (turingtest):

well I suppose the characteristic polynomial for such a linear ODE would have to be\[(r-1)(r-2)(r-3)=0\]multiply that out and that should tell you how to get your ODE

OpenStudy (anonymous):

I get the char. eqn as \[r ^{3}-6r ^{2}+11r-6\] Do I replace the r with respective derivative of y? i.e. \[r ^{2}\] with y" ?

OpenStudy (turingtest):

exactly

OpenStudy (anonymous):

in that case I get:\[y''' - 6y'' + 11y - 6 = 0\] Now if I subtitute y=e^x in this equation, since e^x is a solution, the LHS doesn't come out to 0? Shouldn't it equate to 0?

OpenStudy (turingtest):

careful, where is your y' term?

OpenStudy (anonymous):

Aha right it should be y′′′−6y′′+11y'−6y=0

OpenStudy (turingtest):

right, I hope that works haven't checked myself

OpenStudy (anonymous):

Yeah that works fine. Well thanks a lot!

OpenStudy (turingtest):

welcome :)